# Correlated Noise in Epoch-Based Stochastic Gradient Descent: Implications for Weight Variances

### v2

@@ Appendix I Loss and Accuracy during Training
The exact variance equation of Theorem 4.3 described in Section 4.3 and the calculation shown in Appendix C are to some extent still valid even if the previously mentioned assumption $[{\bf C},{\bf H}]\neq 0$ is not given. In particular, the result for the correlation time holds independently of the commutativity assumption. However, the calculated weight and velocity variances are no longer eigenvalues of the corresponding covariance matrices, but only the variances in the directions of the chosen eigenvector of the Hessian matrix.
@@
Assuming that ${\bf C}$ and ${\bf H}$ do not necessarily commute, we can still project the update equations onto an arbitrary Hessian eigenvector $\bm{\mathbf{p}}_{i}$ with eigenvalue $\lambda_{i}$, which gives us
$v_{k,i}$ $=-\eta\lambda_{i}\theta_{k-1,i}+\beta v_{k,i}-\eta\bm{\mathbf{p}}_{i}\cdot\bm{\mathbf{\delta g}}_{k}\ ,$ (116a)
$\theta_{k,i}$ $=(1-\eta\lambda_{i})\theta_{k-1,i}+\beta v_{k,i}-\eta\bm{\mathbf{p}}_{i}\cdot\bm{\mathbf{\delta g}}_{k}\ .$ (116b)
@@
Since $\bm{\mathbf{p}}_{i}\cdot\bm{\mathbf{\delta g}}_{k}$ is independent of weights and velocity, these two equations are decoupled for each individual Hessian eigenvector. As $\bm{\mathbf{p}}_{i}\cdot\bm{\mathbf{\delta g}}_{k}$ still follows the proposed anti-correlation, the calculations of Appendix C can be performed similarly. Therefore, the theory prediction for the correlation time $\tau_{i}$, defined as the ratio between the weight and the velocity variance in the eigendirection $\bm{\mathbf{p}}_{i}$ of the Hessian, is still valid. However, the weight and the velocity variance in the eigendirection $\bm{\mathbf{p}}_{i}$ are no longer eigenvalues of the covariance matrices $\bf\Sigma$ and $\bf\Sigma_{v}$ if $\bm{\mathbf{p}}_{i}$ is not also an eigenvector of ${\bf C}$.
To better understand the consequences of a situation where ${\bf C}$ and ${\bf H}$ do not commute, we perform an SGD simulation in an artificial quadratic loss landscape where we introduce epoch-based noise with anti-correlations as described in Section 4.1. However, we choose the Hessian matrix $\bf H$ of the loss and the gradient noise covariance matrix $\bf C$ to be non-commuting. The results are shown in Figure 13.
When $\bf H$ and $\bf C$ are approximately proportional to each other, to an extent reported in the literature [31], the results still show good agreement with our theory and there are only small deviations for very small Hessian eigenvalues. If there is no similarity between $\bf H$ and $\bf C$, then the results for the variances no longer follow our theory, but the prediction for the correlation time $\tau_{i}$ still shows good agreement with the numerical data, as expected.
Figure 13: Variances and correlation times of an SGD trajectory in an artificial quadratic potential with epoch-based noise. The quantities are examined in the Hessian eigenbasis and plotted against the corresponding eigenvalue. For the top two panels, the covariance matrix of the noise $\vec{C}$ is equal to the Hessian $\vec{H}$ plus a random matrix $\vec{M}$ of small magnitude ($\vec{M}=\frac{1}{d}\vec{X}\vec{X}^{\top}$ with $\vec{X}\in\mathbb{R}^{d\times d}$ and $X_{ij}\sim\mathcal{N}(0,\,\sigma^{2})$, $\sigma=0.02$, $d=500$). The cosine similarity between $\vec{C}$ and $\vec{H}$ is 0.96 which is a value that has been found empirically in real networks before [31]. Here, our theory for the variances still holds except below a certain small eigenvalue. For the bottom two panels, the Hessian and the covariance matrix are two independent random matrices. There is no longer a relationship between the Hessian eigenvalue and the velocity covariance. However, the theoretical prediction for the correlation time still holds.

### v3

@@ G.7 Relation between Hessian and noise covariance
The exact variance equation of Theorem˜4.3 described in Section˜4.3 and the calculation shown in ˜B are still valid even if the previously mentioned assumption ${\bf C},{\bf H}]\neq 0$ is not given. However, the calculated weight and velocity variances are no longer eigenvalues of the corresponding covariance matrices, but variances in the directions of the chosen eigenvector of the Hessian matrix.
@@
Assuming that $\bf C$} and $\bf H$} do not necessarily commute, we can still project the update equations onto an arbitrary Hessian eigenvector $\boldsymbol{\mathbf{p}}_{i}$ with eigenvalue $\lambda_{i}$, which gives us
$v_{k,i}$ $=-\eta\lambda_{i}\theta_{k-1,i}+\beta v_{k,i}-\eta\boldsymbol{\mathbf{p}}_{i}\cdot\boldsymbol{\mathbf{\delta g}}_{k}\ ,$ (140a)
$\theta_{k,i}$ $=(1-\eta\lambda_{i})\theta_{k-1,i}+\beta v_{k,i}-\eta\boldsymbol{\mathbf{p}}_{i}\cdot\boldsymbol{\mathbf{\delta g}}_{k}\ .$ (140b)
Since we assume that $\boldsymbol{\mathbf{p}}_{i}\cdot\boldsymbol{\mathbf{\delta g}}_{k}$ is independent of weights and velocity, these two equations are decoupled for each individual Hessian eigenvector. As $\boldsymbol{\mathbf{p}}_{i}\cdot\boldsymbol{\mathbf{\delta g}}_{k}$ still follows the proposed anti-correlation, the calculations of ˜B can be performed similarly. Therefore, in a case without commutativity, the theory makes predictions for the variances along the Hessian eigenvectors, $\sigma_{\boldsymbol{\mathbf{\theta}},i}^{2}\coloneqq\boldsymbol{\mathbf{p}}_{i}^{\top}\boldsymbol{\mathbf{\Sigma}}\boldsymbol{\mathbf{p}}_{i}$ and $\sigma_{\boldsymbol{\mathbf{v}},i}^{2}\coloneqq\boldsymbol{\mathbf{p}}_{i}^{\top}\boldsymbol{\mathbf{\Sigma_{v}}}\boldsymbol{\mathbf{p}}_{i}$, depending on the noise variance in the given direction, $\sigma_{\boldsymbol{\mathbf{\delta g}},i}^{2}\coloneqq\boldsymbol{\mathbf{p}}_{i}^{\top}\boldsymbol{\mathbf{C}}\boldsymbol{\mathbf{p}}_{i}$, with $\sigma_{\boldsymbol{\mathbf{\delta g}},i}^{2}=\left\langle\left(\boldsymbol{\mathbf{p}}_{i}\cdot\boldsymbol{\mathbf{\delta g}}_{k}\right)^{2}\right\rangle$. Independent of whether $\bf C$} and $\bf H$} commute, the results imply that the weight covariance, when restricted to the Hessian eigenspace corresponding to eigenvalues smaller than the crossover value, $\lambda_{i}<\lambda_{\rm cross}$, denoted by $\boldsymbol{\mathbf{\Sigma}}_{<}$, is reduced compared to a setup without anti-correlations. To illustrate this effect, we examine the trace of the restricted weight covariance, ${\rm Tr}(\boldsymbol{\mathbf{\Sigma}}_{<})$. The theoretical prediction for its relative reduction is given by
$\frac{{\rm Tr}(\boldsymbol{\mathbf{\Sigma}}_{<,\textrm{with anti-correlations}})}{{\rm Tr}(\boldsymbol{\mathbf{\Sigma}}_{<,\textrm{without anti-correlations}})}\approx\frac{\frac{M\eta}{3(1-\beta)}\sum_{i=i_{\rm cross}}^{d}\lambda_{i}\sigma_{\boldsymbol{\mathbf{\delta g}},i}^{2}}{\sum_{i=i_{\rm cross}}^{d}\sigma_{\boldsymbol{\mathbf{\delta g}},i}^{2}}\ ,$ (141)
where $i_{\rm cross}$ is the index of the largest Hessian eigenvalue $\lambda_{i}$ that is still smaller than the crossover value $\lambda_{\rm cross}$. To develop an intuitive understanding of this result, consider a simplified case where all Hessian eigenvalues satisfying $\lambda_{i}<\lambda_{\rm cross}$ are assumed to be equal, such that $\lambda_{i}=\lambda_{\rm small}$. Under this assumption, the reduction formula Equation˜141 simplifies to $\frac{M\eta\lambda_{\rm small}}{3(1-\beta)}$. Since $\lambda_{\rm small}<\lambda_{\rm cross}\coloneqq\frac{3(1-\beta)}{\eta M}$, this expression is necessarily smaller than one. This confirms that the presence of anti-correlations leads to a systematic reduction in the weight covariance without assuming that $\bf C$} and $\bf H$} commute.
@@
Furthermore, the theory prediction for the correlation time $\tau_{i}$, defined as the ratio between the weight and the velocity variance in the eigendirection $\boldsymbol{\mathbf{p}}_{i}$ of the Hessian, is also valid independently of the commutation relation between $\bf C$} and $\bf H$}. However, the weight and the velocity variance in the eigendirection $\boldsymbol{\mathbf{p}}_{i}$ defined above are no longer eigenvalues of the covariance matrices $\bf\Sigma$ and $\bf\Sigma_{v}$ if $\boldsymbol{\mathbf{p}}_{i}$ is not also an eigenvector of $\bf C$}. Nevertheless, these variances are still a good approximation for the actual eigenvalues, as long as $\bf C$} and $\bf H$} commute approximately.
While empirical investigations show that $\bf C$} and $\bf H$} do not commute exactly, they show a high alignment and suggest that the Hessian eigenbasis is a good approximation for the eigenbasis of the noise covariance (see Section˜3.1). We find such an approximate commutativity as well. We investigated the Hessian and the covariance of the recorded noise in the approximated eigenspace of the Hessian during the analysis period of the LeNet considered in the main text. In Figure˜14 one can see, that the actual 200 largest eigenvalues of the noise covariance align well with the variances in the directions of the 200 Hessian eigenvectors corresponding to the largest Hessian eigenvalues, indicating that the Hessian eigenbasis is very similar to the eigenbasis of the noise covariance.
Figure 14: Noise variance during the analysis period of the LeNet considered in the main text. We analyzed the gradient noise recorded in the subspace of the 5,000 largest Hessian eigenvalues and corresponding eigenvectors ${\bf p}_{i}^{\bf H}$. We calculated the covariance matrix of the recorded noise, $\bf C$}, and its eigenvectors ${\bf p}_{i}^{\bf C}$ in descending order according to their eigenvalues. We then plotted the largest eigenvalues of the noise covariance, which is the noise variance along ${\bf p}_{i}^{\bf C}$, together with the noise variance along ${\bf p}_{i}^{\bf H}$. When $\bf C$} and $\bf H$} commute, there exists a simultaneous eigenbasis for both matrices, and the shown variances would align perfectly. We find that the variances show very good alignment, indicating that both matrices commute approximately.
We also investigated how well the approximation of proportionality between $\bf C$} and $\bf H$} holds. For this we calculated the cosine similarity between both matrices in the approximated Hessian eigenspace, which is the normalized dot product between the flattened matrices. We found a cosine similarity of 0.82, which is similar to the empirical similarities found by Thomas et al. [31] and significantly higher than that of two random low-rank matrices. Additionally, we plotted the noise variances along the approximated Hessian eigenvectors against the corresponding Hessian eigenvalues and found that the two indeed follow an approximate linear relationship (see Figure˜15). So, while an exact proportionality between $\bf C$} and $\bf H$} is not satisfied, it seems to be a very good approximation.
Figure 15: Relationship between Hessian eigenvalues and the variances of the recorded noise during the analysis period of the LeNet considered in the main text. The solid line signifies a linear fit. The exponent resulting from the power law relationship is $1.075\pm 0.002$ with a $2\sigma$-error. If the noise covariance were proportional to the Hessian matrix, the exponent should be equal to one.
G.8 Different Network Architecture
To further confirm our theoretical predictions within trained networks, in this section we turn to a more modern architecture. Instead of the previously used LeNet network, we examine the ResNet-20 network [5]. It is a convolutional network with significantly more convolutional layers than LeNet. It also uses residual blocks with residual connections, which allows for deeper network structures. As our loss function, we again employed Cross Entropy, along with an L2 regularization with a prefactor of $5\cdot 10^{-3}$ and we did not use batch normalization. The number of layers, which is already indicated in the name with 20, is significantly higher than in the LeNet with just five layers. With approximately 272,000 parameters, the ResNet-20 also has significantly more parameters and the computational cost is significantly higher.
Therefore, in this section we limit ourselves to examining the weight and velocity variances in the subspace of the 400 largest Hessian eigenvalues and associated eigenvectors. The network was trained with SGD for 300 epochs using the same exponential learning rate schedule as before, with a learning rate of $5\cdot 10^{-3}$, a momentum parameter of $0.9$, and a minibatch size of $50$. This setup achieves 100% training accuracy and 76% testing accuracy. In Figure˜16 one can observe a good agreement between the theory predictions for the variances and correlation times and the numerical observations.
Figure 16: Relationship between Hessian eigenvalues and the variances and correlation times for a ResNet-20 trained on CIFAR10. In the left panel, we present the variances of weights and velocities. The solid lines signify theoretical predictions from Equation˜6, assuming $\boldsymbol{\mathbf{C}}\approx c_{0}\boldsymbol{\mathbf{H}}$ with $c_{0}$ fitted to the data. The right panel showcases the correlation time together with the theoretical prediction resulting from Equation˜6, which does not require the $\boldsymbol{\mathbf{C}}\approx c_{0}\boldsymbol{\mathbf{H}}$ assumption. The analysis was performed for the 400 largest Hessian eigenvalues and corresponding eigendirections.